# BAP v2: An Enhanced Task Framework for Instruction Following in Minecraft Dialogues

### v2

@@ 2.3 Formalizing the BAP task
In Minecraft, blocks can be placed into the cells $\mathbf{c}=\langle x,y,z\rangle$ of a discrete 3D grid if $\mathbf{c}$ is empty and either on the ground or adjacent to any cell $\mathbf{c}^{\prime}$ that contains a block. A cell $\mathbf{c}=\langle x,y,z\rangle$ is on the ground if its height $y=1$, and is adjacent to any cell $\mathbf{c}^{\prime}=\langle x^{\prime},y^{\prime},z^{\prime}\rangle$ if they touch on one side. [^1] Cells $\langle x,y,z\rangle$ and $\langle x^{\prime},y^{\prime},z^{\prime}\rangle$ are adjacent if either $x=x^{\prime}\pm 1,y=y^{\prime}$, and $z=z^{\prime}$, or $x=x^{\prime},y=y^{\prime}\pm 1$, and $z=z^{\prime}$, or $x=x^{\prime},y=y^{\prime}$, and $z=z^{\prime}\pm 1$ In the MCBT, there are six different BlockColors, and Builders place blocks inside a predefined build region BuildRegion, an $11\times 9\times 11$ box that contains all grid cells $\mathbf{c}=\langle x,y,z\rangle$ from $\langle x_{\min},y_{\min},z_{\min}\rangle$ to $\langle x_{\max},y_{\max},z_{\max}\rangle$. [^2] In the Minecraft Dialogue Corpus, the BuildRegion’s horizontal coordinates $x$ and $z$ range from $x_{\min}=z_{\min}=-5$ to $x_{\max}=z_{\max}=+5$ and vertical coordinates $y$ range from $y_{\min}=1$ to $y_{\max}=9$. The current Built structure $S$ can be represented as a list of blocks and their locations:
@@
The Builder can move freely around the environment. Since spatial relations in the instructions (“the block in front of you", “to your left") often depend on the Builder’s current location and viewpoint, we record the position and pose/orientation of the Builder agent $\textsc{Pos}_{\textbf{B}}=(\langle x_{\textbf{B}},y_{\textbf{B}},z_{\textbf{B}}\rangle,\phi_{\textbf{B}},\gamma_{\textbf{B}})$, where $\langle x_{\textbf{B}},y_{\textbf{B}},z_{\textbf{B}}\rangle$ is a (real-valued) location (that may or may not be inside the build region), $\phi_{\textbf{B}}\in[-90,...,+90]$ indicates a pitch (vertical rotation) and $\gamma_{\textbf{B}}\in[-180,...,+180]$ a yaw (horizontal orientation). Note that $\textsc{Pos}_{\textbf{B}}$ is a critical component of game context as it dictates how spatial relations are interpreted.
@@
[…]
@@ 2.4 Evaluating BAP models
[…]
@@
Strict $\textbf{Strict Precision}~{}P_{\mathrm{strict}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(A_{M}^{\textrm{net}}\cap A_{H}^{\textrm{net}})|}{|A_{M}^{\textrm{net}}|}$
Strict $\textbf{Strict Recall}~{}R_{\mathrm{strict}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(A_{M}^{\textrm{net}}\cap A_{H}^{\textrm{net}})|}{|A_{H}^{\textrm{net}}|}$
[…]
2.5 Data Augmentation
Since the small size of the training set (3,709 examples) is a major limiting factor for training complex models, in Jayannavar, Narayan-Chen, and Hockenmaier (2020), we generated synthetic variants of the original game logs in the training data by combining three data augmentation techniques: utterance paraphrasing (synonym-based substitutions), color substitution (random color permutations across logs), and spatial transformations (rotating structures and Builder B). Empirically, we found that increasing the training data to 14,836 (4x) items gave the best performance for our baseline model (Section 2.6). We will refer to this as the original BAP (augmented) data.
@@ 2.6 A CNN- and GRU-based baseline model
In Jayannavar, Narayan-Chen, and Hockenmaier (2020), we proposed a neural model for the BAP task. In this work, we build upon it and refer to it as the baseline model. The model (Figure 4) is based on a recurrent encoder-decoder architecture (Sutskever, Vinyals, and Le, 2014; Cho et al., 2014) in which a GRU-based encoder (bottom left box) captures the game context (dialogue and action history) via GloVe embeddings (Pennington, Socher, and Manning, 2014), and a CNN-based encoder (top left box) captures the world state at each time step. The world state encompasses the current built structure, action history and the Builder’s position and orientation. The decoder (right box) predicts one action per time step, based on the game history, the world state at that time, and the last action taken. It consists of another GRU backbone over action sequences (bottom right), and a multi-class classifier that reads in the output of the GRU backbone as well as the world state encoding produced by the CNN to predict either the next action (block placement or removal) to be taken, or a special STOP token that terminates the action sequence. Actions are represented as feature vectors and embedded via feedforward layers before being fed to the decoder GRU. The world state representation gets updated and re-encoded after each predicted action. We train this model with teacher forcing and cross entropy loss on the original BAP (augmented) data, and use greedy decoding (max. sequence length of 10 actions) at test time. This yields an F1 score of 21.2% on the BAP test set.
@@ 3 Revisiting BAP Evaluation: A Cleaner Test Set
Since the MDC consists of real human-human game logs collected in a low-stakes, free-flowing setting with minimal constraints on the participants’ language and actions, rather than standalone instructions, this strategy does not always yield BAP items that are suitable for evaluation. First, Builders might move blocks even when no new instructions were given, or before it is completely clear what they should be building next (e.g. when the Architect only instructs the Builder to “build a rectangle to the right of that", without specifying the size or color of the rectangle, let alone its orientation and distance to “that"). In such cases, the corresponding BAP items have an unclear context. Second, Builders occasionally make mistakes, and are frequently interrupted by the Architect, who may want to provide more information or simply affirm that the Builder’s actions are correct. In those cases, the BAP item’s action sequence may lead to an incorrect structure. Finally, when the board is empty, we observe that Architects do not always identify a specific location or orientation for the blocks to be placed, even when they clearly describe the structure to be built (e.g., the instruction “place a red block on the ground"). This likely occurs because there are no blocks on the board or landmarks in the environment that could serve as spatial anchor to be mentioned, and because Architects are aware that the game only requires the Builder’s structure to match the Target as long as it appears anywhere inside the Build Region and in any orientation. In those situations, two different Builders might start the game by placing the same structure in two different locations and/or orientations and both might be correct, since the context provides multiple (valid) interpretations. This phenomenon arises almost exclusively with an empty board, since the blocks on a non-empty board and the current position of the Builder typically provide enough constraints for a clear instruction to only have a unique (valid interpretation) in that case. In this section, we perform an in-depth analysis of the items in the BAP test set that identifies these phenomena, and introduce a new, cleaner BAP test set that should be used to evaluate BAP models instead of the original, legacy test set. In the next section, we revisit the strict F1 metric, and define a fairer variant that accounts for the fact that some empty-board test items have multiple valid interpretations.
3.1 Annotating the test set
We manually annotate items in the test set so that we can identify cases that have to be distinguished for evaluation purposes. Annotators were asked to annotate all items in the same dialog in sequence. For each non-empty board (NEB) item, annotators were first asked whether the context is clear and has a unique interpretation , i.e. clearly specifies what structure to be built next, and there is only a single way of placing that structure on the board (as is typically the case when the board is not empty), or whether the context is unclear, i.e. it is unclear what the Builder should do next, either because more information is required, or no new instruction was given. For empty board (EB) items, annotators could also indicate that the context is clear and has multiple interpretations , i.e. it is clear what structure should be built, even though there are multiple possible placements or orientations within the build region, as is often the case when the board is empty. If the context is clear (with unique or multiple interpretations), a second question asked whether the built structure is correct or incorrect. Details about the annotation setup are provided in the appendix (Section 14). We now illustrate the resulting categories of BAP items with a few examples.
Clear context with unique interpretation and correct structure (Figure 5)
It is clear what structure B should build next, only one such structure can be built, and the human Builder actually built that structure (Figure 5(a)). This can also arise when the board is empty (Figure 5(b)), since B is first made to stand at the bottom right corner of the board, and A clearly specifies the locations of the two blocks to be placed, given B’s new orientation.
Clear context with unique interpretation and incorrect structure (Figure 6)
Here, it is clear what structure B should build next (and only one such structure can be built), but B’s actions did not result in that structure, perhaps because B got interrupted by A, or because B made a genuine mistake. In Figure 6(a), A wants B to place two red blocks directly below the middle block of the top of the arch, but B only places one of those blocks. In Figure 6(b), A wants B to place a red block diagonally above (“on top and in front of") the topmost green block, but B places a red block in the wrong location. In the next BAP item from that dialog (Figure 6(c)), A tries to correct B’s mistake, but B now places two red blocks, only one of which should be there in the final structure.
Clear context with multiple interpretations and correct structure (Figure 7)
When the board is empty, it is often clear what structure B should build next, but the placement and orientation of that structure may not have been uniquely specified. In this case, the context is clear, although it can be (correctly) interpreted in multiple ways.
Unclear context (Figure 8)
Finally, some action sequences occur in contexts in which it is unclear what the Builder should do next. In Figure 8(a), it is unclear what exact location A implies by “the side of the red". In Figure 8(b), A only signals they had made an error, but it is unclear what the error is or what its recency is and A does not provide any information that would allow B to correct the error.
((a)) Item 1
((b)) Item 2
Figure 5: Clear context with unique interpretation and correct structure
((a)) Item 1
((b)) Item 2
((c)) Item 3
Figure 6: Clear context with unique interpretation and incorrect structure
@@ 3.1 Annotating the test set
Figure 7: Empty board: Clear context with multiple interpretations and correct structure
((a)) Item 1
((b)) Item 2
Figure 8: Unclear context
3.2 Creating a clean test set
To obtain a clean BAP test set, we remove all items with unclear contexts, as the the correctness of the built structure cannot (and should not) be assessed for the BAP task. We also automatically fix/replace incorrect structures that can be fixed in such manner, and remove the items with structures that can’t be fixed. This results in a new, BAP v2 test set that contains only items with clear contexts and correct structures, and thus suitable for fairer BAP evaluation.
Fixing incorrect structures by looking ahead to the next correct structure
If the Builder made a mistake or their action sequence was interrupted by an utterance, the built structure of a BAP item may be incorrect. E.g. in Figure 9(a) B places only one instead of two blocks. However, the next BAP item (Figure 9(b)) has a correct structure, and its structure is the one that should have been built in 9(a) because 9(b) completes the truncated action sequence. We can therefore replace the structure in 9(a) with the one in 9(b). Figure 10 shows a sequence of three BAP items: in 10(a) B misinterprets A and makes an error, in 10(b) B realizes the mistake post A’s correction and proceeds to correct it, but A interrupts them, and in 10(c), the mistake is finally rectified post the interruption. Items 10(a) and 10(b) with incorrect structures should therefore be amended to result in the correct structure built in 10(c).
Manual analysis shows that sequences of one or more incorrect items (with clear contexts) that are immediately followed by a correct item (with clear context) can be fixed by replacing their built structures with the built structure of that subsequent correct item. Those items with incorrect structures that conform to this pattern are thus fixed automatically, and the remaining that can’t are excluded from the final test set.
((a)) Item $i$ has an incorrect built structure
((b)) Item $i+1$ has a correct built structure
Figure 9: An item with an incorrect built structure (9(a)) that can be fixed by the correct built structure in the following item (9(b))
((a)) Item $i$ has an incorrect built structure
((b)) Item $i+1$ has an incorrect built structure
((c)) Item $i+2$ has the correct built structure for items $i$, $i+1$ and $i+2$
@@ 3.2 Creating a clean test set
Figure 10: A sequence of items with incorrect built structures (10(a), 10(b)) that can be fixed by the following item’s correct built structure (10(c))
Statistics of the clean BAP v2 test set

### v3

@@ 2.4 Formalizing the BAP task
In Minecraft, blocks can be placed into the cells $\mathbf{c}=\langle x,y,z\rangle$ of a discrete 3D grid if $\mathbf{c}$ is empty and either on the ground or adjacent to any cell $\mathbf{c}^{\prime}$ that contains a block. A cell $\mathbf{c}=\langle x,y,z\rangle$ is on the ground if its height $y=1$, and is adjacent to any cell $\mathbf{c}^{\prime}=\langle x^{\prime},y^{\prime},z^{\prime}\rangle$ if they touch on one side. [^2] Cells $\langle x,y,z\rangle$ and $\langle x^{\prime},y^{\prime},z^{\prime}\rangle$ are adjacent if either $x=x^{\prime}\pm 1,y=y^{\prime}$, and $z=z^{\prime}$, or $x=x^{\prime},y=y^{\prime}\pm 1$, and $z=z^{\prime}$, or $x=x^{\prime},y=y^{\prime}$, and $z=z^{\prime}\pm 1$ In the MCBT, there are six different BlockColors, and Builders place blocks inside a predefined build region BuildRegion, an $11\times 9\times 11$ box that contains all grid cells $\mathbf{c}=\langle x,y,z\rangle$ from $\langle x_{\min},y_{\min},z_{\min}\rangle$ to $\langle x_{\max},y_{\max},z_{\max}\rangle$. [^3] In the Minecraft Dialogue Corpus, the BuildRegion’s horizontal coordinates $x$ and $z$ range from $x_{\min}=z_{\min}=-5$ to $x_{\max}=z_{\max}=+5$ and vertical coordinates $y$ range from $y_{\min}=1$ to $y_{\max}=9$.
The current Built structure $S$ can be represented as a list of blocks and their locations:
@@
The Builder can move freely around the environment. Since spatial relations in the instructions (“the block in front of you", “to your left") often depend on the Builder’s current location and viewpoint, we record the position and pose/orientation of the Builder agent $\textsc{Pos}_{\textbf{B}}=(\langle x_{\textbf{B}},y_{\textbf{B}},z_{\textbf{B}}\rangle,\phi_{\textbf{B}},\gamma_{\textbf{B}})$, where $\langle x_{\textbf{B}},y_{\textbf{B}},z_{\textbf{B}}\rangle$ is a (real-valued) location (that may or may not be inside the build region), $\phi_{\textbf{B}}\in[-90,...,+90]$ indicates a pitch (vertical rotation) and $\gamma_{\textbf{B}}\in[-180,...,+180]$ a yaw (horizontal orientation) angle. For brevity, we use $B^{t}$ to refer to $\textsc{Pos}_{\textbf{B}}$ at time $t$ in the formal definition of BAP items.
@@
We specify Builder actions $a=(t,c,\mathbf{c})$ as triplets consisting of an action type $t\in\mathcal{T}=\{\textsc{Place},\textsc{Remove}\}$, a block color $c\in\textsc{BlockColors}$, and a grid cell $\mathbf{c}=\langle x,y,z\rangle\in\textsc{BuildRegion}$ whose state is changed by aa. [^4] Since the Malmo API can be used to move agents into positions in which a particular block is accessible to them, we ignore the Builder’s movement, and only consider its block placement and removal actions. If a feasible action $a$ is executed, it changes the built structure $S$ to $S^{\prime}\neq S$: $S~\xrightarrow{a}~S^{\prime}$. A placement is feasible if $\mathbf{c}$ is empty and either on the ground or adjacent to a non-empty cell, and results the addition of a $c$-colored block located in $\mathbf{c}$ to $S^{\prime}$. A removal action is feasible if $S$ contains a contains a block of color $c$ in $\mathbf{c}$, and results the exclusion of that block from $S^{\prime}$. When Builders follow an instruction, they execute an action sequence $A=\langle a^{1},...,a^{k}\rangle$ for $k\geq 1$ and $a^{1},...,a^{k}\in\textsc{ActionSpace}$ that changes $S$ to $S^{\prime}$: $S\xrightarrow{A=\langle a^{1},...a^{k}\rangle}S^{\prime}$. ActionSpace is the set of all 7623 possible actions (7 actions available per cell in the $11\times 9\times 11$ build region—placing a block in one of 6 colors or removing a block).
@@ 2.5 Evaluating BAP models
To compare two action sequences, or to measure the change from $S$ to $S^{\prime}$, it is helpful to note that any action $a=(t,c,\mathbf{c})$ is undone by the inverse action $a^{-}=(t^{-},c,\mathbf{c})$ where $t^{-}=\textsc{Remove}$ if $t=\textsc{Place}$, and $t^{-}=\textsc{Place}$ if $t=\textsc{Remove}$. Most commonly, this occurs when blocks are placed to serve as a necessary supporting block for a floating structure and removed in the same action sequence. These supporting blocks can be ignored since they do not occur in $S^{\prime}$. Moreover, builders are free to use any color they wish, and can often choose among a variety of possible locations where to place these blocks. Human Builders are also prone to accidentally placing or removing blocks, but typically recognize and correct such mistakes immediately within the same action sequence. Finally, many structures do not require their blocks to be placed in a particular order. We can therefore transform any (feasible) action sequence $A$ into a set of net actions $A^{\mathrm{net}}$ by first removing from $A$ any actions $a$ and their inverse $a^{-}$ and turning the resulting (infeasible) sequence $A^{\prime}$ into a set $A^{\mathrm{net}}$. The distance $\Delta(S,S^{\prime})$ between two structures $S$ and $S^{\prime}$ can then be defined as the number of net actions of any action sequence $A:S\xrightarrow{A}S^{\prime}$ that changes $S$ to $S^{\prime}$: $\Delta(S,S^{\prime})=|A^{\mathrm{net}}|$, and any two action sequences $A_{1}$ and $A_{2}$ are equivalent ($A_{1}\equiv A_{2}$), i.e. lead to the same subsequent structure, if they have the same set of net actions ($A_{1}^{\mathrm{net}}=A_{2}^{\mathrm{net}}$). We therefore evaluate BAP models against human Builders by comparing their respective net actions, $A_{M}^{\textrm{net}}$ and $A_{H}^{\textrm{net}}$ to compute a (strict) F1 score (we report a micro-averaged F1 score over all action sequences in the test/dev data):
@@
$\textbf{Strict Precision}~P_{\mathrm{strict}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(A_{M}^{\textrm{net}}\cap A_{H}^{\textrm{net}})|}{|A_{M}^{\textrm{net}}|}$
$\textbf{Strict Recall}~R_{\mathrm{strict}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(A_{M}^{\textrm{net}}\cap A_{H}^{\textrm{net}})|}{|A_{H}^{\textrm{net}}|}$
$\textbf{Strict F1}~F1_{\mathrm{strict}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{2\cdot P_{\mathrm{strict}}\cdot R_{\mathrm{strict}}}{P_{\mathrm{strict}}+R_{\mathrm{strict}}}$
@@ 2.6 A CNN- and GRU-based baseline model
In Jayannavar, Narayan-Chen, and Hockenmaier (2020), we proposed a neural model for the BAP task. In this work, we build upon it and refer to it as the GRU-based baseline model. The model is based on a recurrent encoder-decoder architecture (Sutskever, Vinyals, and Le, 2014; Cho et al., 2014) in which a GRU-based encoder captures the game context (dialogue and action history) via GloVe embeddings (Pennington, Socher, and Manning, 2014), and a CNN-based encoder captures the world state at each time step. The world state encompasses the current built structure, action history and the Builder’s position and orientation. We trained this model with teacher forcing and cross entropy loss on the original BAP data, and used greedy decoding (max. sequence length of 10 actions) at test time. This yields an F1 score of 21.2% on the BAP test set.
2.7 Related Work: LLMs for BAP
Related to our work are the recent works of Chaturvedi, Thompson, and Asher (2024) and Kranti, Hakimov, and Schlangen (2024), both exploring more modern LLMs for the BAP task. The former finetunes the Llama-3-8B model (Dubey et al., 2024) using QLoRA (Dettmers et al., 2023) on the original BAP data, resulting in the Nebula model, which predicts the next action sequence in a text-to-text fashion using all preceding conversation and action sequences as context. The latter employs in-context learning (ICL) with GPT-4. Both achieve state-of-the-art results over the GRU-based model. Nebula achieves an F1 score of 39.2% and GPT-4 scores 39.0% (under the now legacy BAP evaluation), demonstrating that both approaches yield very similar top-line performance (Nebula being marginally better). A slightly more detailed comparison to Chaturvedi, Thompson, and Asher (2024) is provided later in Section 7.4 given some other relevant aspects of their work.
2.8 Related Work: Tasks and Datasets
In this section, we explain why the BAP task continues to be a relevant and challenging testbed for grounded instruction following. The MDC dataset is a collaborative embodied task, while BAP focuses on grounded language instruction following. General collaborative task-oriented dialogue tends to focus on un-embodied domains. Typical of the domain is the MultiWoz dataset (Budzianowski et al., 2018) which explores conversations around tasks such as trip planning, which require both dialogue and information retrieval. Even tasks that include a visual modality are not explicitly embodied: take Udagawa and Aizawa (2019)’s OneCommon task, which requires two agents to work together to identify the overlap in their fields of view, yet does not include turning to adjust a perspective, etc.
The MDC, on the other hand, is closer to some vision-and-language works. Vision-and-Language Navigation (VLN) (Anderson et al., 2018), and its dialogue counterpart, Cooperative Vision-and-Dialog Navigation (CVDN) (Thomason et al., 2019), which explore instruction-following in photorealistic navigation settings, are good examples. Within this domain, others have combined vision and dialogue navigation with task completion, including CEREALBAR (Suhr et al., 2019), a game task in which an instruction-follower and an instruction-giver cooperate to traverse a map and collect items. Others have combined vision and language-based task completion in ways that go beyond navigation. Roy et al. (2021) consider a human and a robot collaborating to search for an object, while ALFRED (Shridhar et al., 2020) is a benchmark for mapping natural language instructions and egocentric vision to action sequences for household tasks in simulated environments. Gao et al. (2022) extend this by incorporating dialogue instead of just instructions. Many of the challenging aspects of the MDC are shared across these tasks, e.g. asking for clarification or help is a central component in VLN works Nguyen and Daumé (2019); Chi et al. (2019). Central to all of these tasks is that agents and humans must communicate with natural language about a world that exists outside of text. Although more recent tasks require real vision to represent this world, their underlying world state space (defined by fixed viewpoints and underlying navigation graphs) is just as highly discretized as a Minecraft grid. And, while many works do opt for visual inputs, Jia et al. (2024) explore the potential of leaving vision out of task-oriented dialogue, and works in less grounded settings, such as Xu et al. (2024), often utilize fully textual agents. Despite not requiring vision as most of the counterparts discussed above do, our task poses a possibly more challenging planning problem: its action space is much larger (7623 possible actions vs. six actions in the vision-language navigation work). By excluding vision, MDC can focus solely on instruction following and the decision-making inherent in that. Furthermore, Minecraft dialogues are more complex than the other discussed datasets due to the MDC’s asynchronous communication, longer turn sequences, and dynamic environment changes during construction. Unlike navigation or simple cooperative games, where referring expressions target existing objects, construction often involves instructions for objects yet to be built.
The TEACh benchmark (Padmakumar et al., 2022) addresses the former of these limitations with free-form dialogue, similar to the MDC, but not the latter as the task does not require references to unbuilt objects. While it does feature object state changes in its 3000+ human-human dialogues, the instruction giver also receives a predefined plan of steps, limiting the diversity since there is only one way to proceed (compared to the many unique ways to build a final MDC structure). Additionally, the MDC allows for greater abstraction in instructions, such as referring to substructures, individual blocks, or shapes, whereas TEACh focuses on well-defined objects in the environment. This combination of dialogue structure, references to unbuilt objects, plan diversity, and the instruction abstraction make the MDC and BAP tasks both difficult and compelling, as well as providing areas of exploration not addressed by other similar tasks.
Minecraft Environment
Among Minecraft-related efforts, IGLU is most closely related to the MDC/MCBT and was directly inspired by it. Unlike the full dialogue structure of the MDC, the IGLU dataset contains mostly single turns (i.e. instruction-action pairs without a broader game context or history). A small multi-turn subset is available (127 dialogues), but has strict turn taking in place of MDC’s free-form asynchronous dialogue. It further simplifies the task by using cardinal directions (e.g “North", “South") which are absolute and more consistent that relative references (e.g. “left", “right").
Some other works also build on or are inspired by the MCBT/MDC: Bonn et al. (2020), Thompson, Hunter, and Asher (2024), Bonial et al. (2021), and Madge et al. (2025) provide additional linguistic annotations for the MDC, while Shi, Feng, and Lipani (2022) explore when the Builder should ask for clarification or execute an instruction, although their building task consists of predicting single actions instead of action sequences, and is therefore not directly comparable to ours.
@@ 3 Revisiting BAP Evaluation: A Cleaner Test Set
Suitable BAP test items (with correct structures)
Clear context with unique interpretation (non-empty board)
Clear context with unique interpretation (empty board)
Clear context with multiple interpretations (empty board)
@@
Figure 4: BAP items with clear contexts and correct structures are suitable for evaluation purposes. If the board is empty, there may be multiple valid interpretations.
Unsuitable BAP test items: Unclear contexts
Figure 5: BAP items with unclear contexts are unsuitable for evaluation purposes
Unsuitable BAP test items: Incorrect structures
Figure 6: BAP items with incorrect structures are unsuitable for evaluation purposes.
Fixable BAP items with incorrect structures
The structure of item $i$ can be replaced by the structure of item $i+1$
Item $i$
Item $i+1$
The structures of item $i$ and $i+1$ can be replaced by the structure of item $i+2$
Item $i$
Item $i+1$
Item $i+2$
@@
Figure 7: A correct structure following one or more BAP items with incorrect built structures can be used to fix the incorrect BAP items.
Since the MDC consists of real human-human game logs collected in a low-stakes, free-flowing setting with minimal constraints on the participants’ language and actions, rather than standalone instructions, not all action sequences correspond to BAP items that are suitable for evaluation purposes. First, Builders might move blocks even when no new instructions were given, or before it is completely clear what they should be building next (e.g. when the Architect only instructs the Builder to “build a rectangle to the right of that", without specifying the size or color of the rectangle, let alone its orientation and distance to “that"). In such cases, the corresponding BAP items have an unclear context. Second, Builders occasionally make mistakes, and are frequently interrupted by the Architect, who may want to provide more information or simply affirm that the Builder’s actions are correct. In those cases, the BAP item’s action sequence may lead to an incorrect structure. Finally, when the board is empty, we observe that Architects do not always identify a specific location or orientation for the blocks to be placed, even when they clearly describe the structure to be built (e.g., the instruction “place a red block on the ground"). This likely occurs because there are no blocks on the board or landmarks in the environment that could serve as spatial anchor to be mentioned, and because Architects are aware that the game only requires the Builder’s structure to match the Target as long as it appears anywhere inside the Build Region and in any orientation. In those situations, two different Builders might start the game by placing the same structure in two different locations and/or orientations and both might be correct, since the context provides multiple (valid) interpretations. This phenomenon arises almost exclusively with an empty board, since the blocks on a non-empty board and the current position of the Builder typically provide enough constraints for a clear instruction to only have a unique (valid interpretation) in that case. In this section, we perform an in-depth analysis of the items in the BAP test set that identifies these phenomena, and introduce a new, cleaner BAP test set that should be used to evaluate BAP models instead of the original, legacy test set. In the next section, we revisit the strict F1 metric, and define a fairer variant that accounts for the fact that some empty-board test items have multiple valid interpretations.
3.1 Identifying suitable BAP test items
To distinguish the cases mentioned above, we manually annotate all BAP items in the test set (for details, see Appendix Section 12). For each non-empty board (NEB) item, annotators were asked whether the context is clear and has a unique interpretation (i.e. it is clear what structure to be built next, and where to place it on the board), or whether the context is unclear (i.e. it is unclear what to build next, e.g. because more information is required, or no new instruction was given). For empty board (EB) items, annotators could also indicate that the context is clear and has multiple interpretations, (i.e. it is clear what structure should be built, even though there are multiple possible placements or orientations of that structure on the empty board.) If the context is clear, a second question asked whether the built structure is correct or incorrect.
Suitable BAP test items
Items whose contexts are clear and whose structures are correct can clearly be used for evaluation purposes (Figure 4). If the board is non-empty, this requires contexts to have a unique interpretation, but if the board is empty, Builders can chose any of typically many correct interpretations.
Unsuitable BAP test items
Items are unsuitable for evaluation if the context is unclear (e.g. because the Architect gave unclear instructions, Figure 5) or if the structure is incorrect (e.g. the Builder made mistakes, Figure 6).
Fixable BAP test items
Manual analysis shows that sequences of one or more incorrect built structures that are immediately followed by a correct structure can be fixed by replacing the incorrect structures with the built structure of that subsequent correct item (Figure 7). Such cases occur typically because the Builder’s (correct) action sequence was interrupted by an utterance, or because the Architect issued a quick correction that the Builder immediately executed.
3.2 Statistics of the clean BAP v2 test set